OpenStudy (anonymous):

If f(x)=-3x-1, find f(4)

myininaya (myininaya):

just replace all the x's in f(x)=-3x-1 with 4

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

If f(x)=-x2+x divide by x, find f(2)

myininaya (myininaya):

@spandana please don't just give the answer Please try to get the asker involved next time. Thanks.

OpenStudy (anonymous):

Do you know the answer?

myininaya (myininaya):

you should try this one @BethMoses2008 f(2) means to replace all the x's with 2's in f(x)=-x^2+x

OpenStudy (anonymous):

okay thanks

myininaya (myininaya):

what did you get when you replaced all the x's with 2's?

OpenStudy (anonymous):

@BethMoses2008 Since you know how to do it, can you find the answer?

OpenStudy (anonymous):

Is 4

OpenStudy (anonymous):

4^3+1=65

myininaya (myininaya):

ok first step i asked you to do what to replace all the x's with 2's so you have -2^2+2 what is 2^2?

OpenStudy (anonymous):

If f(1) f(x)=4x^3+x, find f(1)

OpenStudy (anonymous):

I think it's 65

myininaya (myininaya):

I'm trying to tell you answered incorrectly to the previous one the answer is not 4.

myininaya (myininaya):

-2^2+2 is not 4

OpenStudy (anonymous):

It's 3

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

Typ

OpenStudy (anonymous):

I forgot to add 2

OpenStudy (anonymous):

Nevermind

OpenStudy (anonymous):

If f(x)=-2, find f(5)

myininaya (myininaya):

well 3 is not the right answer to evaluating -2^2+2 let's try to take it slow you are to apply order of operations there are no grouping symbols so let's do the exponents the exponent is only on the 2 if were to be on the -2 it would look like (-2)^2 but we have -2^2 so first exponents what is 2^2?

OpenStudy (anonymous):

If f(x)=-2, find f(5)

myininaya (myininaya):

if you aren't willing to answer my questions, then i cannot help you. sorry.